OpenStudy (anonymous):

Which is equivalent to 6/9 A. 18/36 B. 1/3 C. 3/4 D. 12/18

OpenStudy (anonymous):

I think it's A

OpenStudy (anonymous):

Do me a favor, simplify 6/9 to the smallest fraction you can.

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

No, divide the numerator by 3, and the denominator by 3...

OpenStudy (anonymous):

6/3 ---- =? 9/3

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

Good, now lets simplify 18/36... Divide the numerator AND the denominator by 9.

OpenStudy (anonymous):

OOPS that was wrong. divide both by 18.

OpenStudy (timaashorty):

18/36 divide them with the greastest multiple of the denominator and numerator which is 18. 18/ 18 = 1 ___ ____ 36 / 18 = 2 So it's not A. Lets try D. Simplify it. dived both sides with the GCM which is 6. 12 / 6 = 2 18/ 6 = 3 So it's D. 12/18 equals to 2/3

OpenStudy (timaashorty):

Sorry I took to long to reply my internet is running slow today.

OpenStudy (timaashorty):

Wait. I didn't explain completely. sorry. we're trying to make it equal to 6/9 If you simplify 6/9 it will equal to 2/3 also.

OpenStudy (anonymous):

Makes sense..a little

OpenStudy (timaashorty):

Do you know how to simplify?

OpenStudy (anonymous):

D.12/18

OpenStudy (anonymous):

Yes,you divide it by 2 in simpliest form.Well I sort of know how to do it. When my teacher taught this like 2 or 3 years ago the kids were being wild so i'm not that good with it..Or I guess you could say im not good with math at all

OpenStudy (timaashorty):

I'll give you an example: \[\frac{ 6 }{ 10 }\] FIRST STEP, FIND ALL THE MULTIPLES OF BOTH NUMBERS. The multiples are the values that can multilply to equal that number. > 6: 1, 2, 3, 6 10, 1,2, 5,10 Now look at the numbers. What is the biggest/greatest multiple that BOTH the 6 and 10 have? 2 is. right? So now divide both numbers by two to simplify it. \[6 \div 2 = 3\] \[10 \div 2 = 5\] So \[6/10 simplified = 3/5\]

OpenStudy (anonymous):

Well,thanks for showing me how to simplify correctly and for showing me,well teaching me how to solve this problem

OpenStudy (timaashorty):

You're very welcome [: